OpenStudy (anonymous):

so revenue between 1987 and 1989 is modeled by f(x)=97(1.5)^x millions of $ x years after 1987. how do i find rate of revenue and actual revenue for 1990?

OpenStudy (anonymous):

*rate of change of revenue

OpenStudy (anonymous):

1990 is three years after 1987. So, f(x)=97(1.5^3) 327.37500 million Rate of change would be derivative with respect to X.

OpenStudy (anonymous):

I see! but then what would the percent rate of change be?

OpenStudy (anonymous):

the rate of change would be d/dx = 97*x*1.5^x-1. then I believe you take the value where x=3 and divide it by the revenue at that point (already given above)